# Two-Edge Connectivity via Pac-Man Gluing

### v1

@@ A.2 Helper Lemmas
Let $H\subseteq E$ be such that $H$ is connected and contains a spanning subgraph of $G$, and let $H_{i}\coloneq H\cap E(G_{i})$ for $i\in\{1,2\}$. If for all $i\in\{1,2\}$ every edge $e\in H_{i}$ is part of a 2EC component of $H_{i}$ or lies on an $x-y$ path in $H_{i}$ where $x,y\in\{u,v,w\}$ such that $x\neq y$ and there exists an $x-y$ path in $H_{j}$ where $j=1$ if $i=2$ and $j=2$ if $i=1$, then $H$ is a 2EC spanning subgraph of $G$.
@@
We show that every edge $e\in H$ is part of a 2EC component of $H$, which proves the lemma because $H$ is connected and spanning. If for all $i\in\{1,2\}$ every edge $e\in H_{i}$ is part of a 2EC component of $H_{i}$, it is also part of a 2EC component of $H$. Otherwise, for all $i\in\{1,2\}$, if $e$ lies on an $x-y$ path in $H_{i}$ where $x,y\in\{u,v,w\}$ such that $x\neq y$ and there exists an $x-y$ path in $H_{j}$ where $j=1$ if $i=2$ and $j=2$ if $i=1$, then $e$ is on a cycle of $H$ as both paths are edge-disjoint. ∎
@@
If the condition in Algorithm 2 in Algorithm 2 applies and both, $H^{\prime}_{1}$ and $H^{\prime}_{2}$ are 2EC spanning subgraphs of $G^{\prime}_{1}$ and $G^{\prime}_{2}$, respectively, then there exist edges $F^{\prime}$ with $|F^{\prime}|\leq 4$ such that $H^{\prime}$ is a 2EC spanning subgraph of $G$.
@@
Let $i\in\{1,2\}$, and let $H^{\prime}_{i}(u)$, $H^{\prime}_{i}(v)$, and $H^{\prime}_{i}(w)$ be the connected components of $H^{\prime}_{i}$ after decontracting $\{u,v,w\}$ that contain $u$, $v$ and $w$, respectively. By Lemma 24, there exist edges $F_{i}\subseteq E(G_{i})$ such that $|F_{i}|\leq 2$ and $H^{\prime}_{i}\cup F_{i}$ is connected in $G_{i}$. Since $H^{\prime}_{i}$ is spanning in $G_{i}$, setting $F\coloneq F_{1}\cup F_{2}$ implies that $|F|\leq 4$ and that $H^{\prime}=H^{\prime}_{1}\cup H^{\prime}_{2}\cup F$ is spanning and connected in $G$. Moreover, observe that for $i\in\{1,2\}$ each edge $e\in H^{\prime}_{i}\cup F_{i}$ is either in a 2EC component of $H^{\prime}_{i}\cup F_{i}$ or lies on an $x-y$ path in $H^{\prime}_{i}\cup F_{i}$, where $x,y\in\{u,v,w\}$, $x\neq y$. Thus, Lemma 25 implies that $H^{\prime}$ is a 2EC spanning subgraph of $G$. ∎
@@
If the condition in Algorithm 2 in Algorithm 2 applies and $\mathsf{RED}(G_{2}^{\mathsf{B1}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{B1}}$, then there exist edges $F^{\mathsf{B1}}$ with $|F^{\mathsf{B1}}|\leq 1$ such that $H^{\mathsf{B1}}$ is a 2EC spanning subgraph of $G$.
@@
Assume w.l.o.g. that $u$ and $v$ belong to the same 2EC component of $\mathsf{OPT}_{1}^{\mathsf{B1}}$. Let $H_{1}^{\mathsf{B1}}\coloneq\mathsf{OPT}_{1}^{\mathsf{B1}}$ and $C_{i}^{\mathsf{B1}}(x)$ be the 2EC component of $H_{i}^{\mathsf{B1}}$ that contains $x$, where $x\in\{u,v,w\}$ and $i\in\{1,2\}$. Since $C_{1}^{\mathsf{B1}}(u)=C_{1}^{\mathsf{B1}}(v)$, we have that in $\mathsf{OPT}_{1}^{\mathsf{B1}}\cup H_{2}^{\mathsf{B1}}$, the vertices of $C_{1}^{\mathsf{B1}}(u)$, $C_{2}^{\mathsf{B1}}(u)$, and $C_{2}^{\mathsf{B1}}(v)$ are in the same 2EC component. Furthermore, in $\mathsf{OPT}_{1}^{\mathsf{B1}}$, there is a connection between $C_{1}^{\mathsf{B1}}(u)$ and $C_{1}^{\mathsf{B1}}(w)$. If $w$ and $v$ are already connected in $H_{2}^{\mathsf{B1}}$ then we can set $F^{\mathsf{B1}}\coloneq\emptyset$ :- ∅ and see that this satisfies the lemma, since $H^{\mathsf{B1}}$ contains a spanning tree and every edge in $H^{\mathsf{B1}}$ is either in the 2EC component containing $u$ and $v$, in a 2EC component containing $w$ or on a $w-v$ path in $G_{i}$, $i\in\{1,2\}$. Otherwise, if $w$ and $v$ are not in the same connected component in $H_{2}^{\mathsf{B1}}$, by Lemma 24, there must be an edge $e$ from $C_{2}^{\mathsf{B1}}(w)$ to either $C_{2}^{\mathsf{B1}}(v)$ or $C_{2}^{\mathsf{B1}}(u)$ in $E(G_{2})\setminus H^{\mathsf{B1}}$. Now it can be easily verified that $F^{\mathsf{B1}}\coloneq\{e\}$ satisfies the lemma, since $H^{\mathsf{B1}}$ contains a spanning tree and every edge in $H^{\mathsf{B1}}$ is either in the 2EC component containing $u$ and $v$, in a 2EC component containing $w$ or on a $w-v$ path in $G_{i}$, $i\in\{1,2\}$. ∎
@@
If the condition in Algorithm 2 in Algorithm 2 applies and $\mathsf{RED}(G_{2}^{\mathsf{B2}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{B2}}$, then there exist edges $F^{\mathsf{B2}}$ with $|F^{\mathsf{B2}}|\leq 2$ such that $H^{\mathsf{B2}}$ is a 2EC spanning subgraph of $G$.
@@
The proof is similar to the previous one. Assume w.l.o.g. that $u$ and $v$ belong to the same 2EC component of $\mathsf{OPT}_{1}^{\mathsf{B2}}$. Let $H_{1}^{\mathsf{B2}}\coloneq\mathsf{OPT}_{1}^{\mathsf{B2}}$ and $C_{i}^{\mathsf{B2}}(x)$ be the 2EC component of $H_{i}^{\mathsf{B2}}$ that contains $x$, where $x\in\{u,v,w\}$ and $i\in\{1,2\}$. Since $C_{1}^{\mathsf{B2}}(u)=C_{1}^{\mathsf{B2}}(v)$, we have that in $\mathsf{OPT}_{1}^{\mathsf{B2}}\cup H_{2}^{\mathsf{B2}}$, the vertices of $C_{1}^{\mathsf{B2}}(u)$, $C_{2}^{\mathsf{B2}}(u)$ and $C_{2}^{\mathsf{B2}}(v)$ are in the same 2EC component. Furthermore, in $\mathsf{OPT}_{1}^{\mathsf{B2}}$, there is no connection between $C_{1}^{\mathsf{B1}}(u)$ and $C_{1}^{\mathsf{B2}}(w)$, since it is of type $\mathsf{B2}$. Hence, by Lemma 24, there must be an edge $e_{1}\in E(G_{1})\setminus\mathsf{OPT}_{1}^{\mathsf{B2}}$. Note that $\mathsf{OPT}_{1}^{\mathsf{B2}}\cup\{e_{1}\}$ is now a solution of type $\mathsf{B1}$ for $G_{1}$, and hence we can do the same as in the proof of the previous lemma and observe that $\lvert F^{\mathsf{B2}}\rvert\leq 2$. ∎
@@
If the condition in Algorithm 2 in Algorithm 2 applies and $\mathsf{RED}(G_{2}^{\mathsf{C1}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C1}}$, then there exist edges $F^{\mathsf{C1}}$ with $|F^{\mathsf{C1}}|\leq 2$ such that $H^{\mathsf{C1}}$ is a 2EC spanning subgraph of $G$.
@@
Let $C_{2}^{\mathsf{C1}}(x)$ be the 2EC component of $H_{2}^{\mathsf{C1}}$ that contains $x$, for $x\in\{u,v,w\}$. Note that $\mathsf{OPT}_{1}^{\mathsf{C1}}$ is connected and spanning for $G_{1}$ and that each edge $e\in\mathsf{OPT}_{1}^{\mathsf{C1}}$ is either in some 2EC component or it is on a $y-z$-path in $\mathsf{OPT}_{1}^{\mathsf{C1}}$ for $y,z\in\{u,v,w\}$, $y\neq z$. We consider three cases. If $H_{2}^{\mathsf{C1}}$ is connected, then we claim that $F^{\mathsf{C1}}\coloneq\emptyset$ :- ∅ satisfies the lemma. In this case every edge of $H_{2}^{\mathsf{C1}}$ is part of a 2EC component of $H_{2}^{\mathsf{C1}}$ or it is on a $y-z$-path in $H_{2}^{\mathsf{C1}}$ for $y,z\in\{u,v,w\}$, $y\neq z$. Hence, Lemma 25 implies that $H_{2}^{\mathsf{C1}}\cup\mathsf{OPT}_{1}^{\mathsf{C1}}$ is a 2EC spanning subgraph of $G$.
Next, assume that $H_{2}^{\mathsf{C1}}$ is composed of two connected components, say w.l.o.g. one containing $u$ and $v$ and the other containing $w$, and observe that each component is 2EC. Then Lemma 24 guarantees the existence of an edge $e$ in $G_{2}$ between $C_{2}^{\mathsf{C1}}(w)$ and $C_{2}^{\mathsf{C1}}(u)$ such that $H_{2}^{\mathsf{C1}}\cup\{e\}$ is connected. Thus, we can analogously to the previous case show that $\mathsf{OPT}_{1}^{\mathsf{C1}}\cup H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ with $F^{\mathsf{C1}}\coloneq\{e\}$ is a 2EC spanning subgraph of $G$, since in $H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ each edge is either in some 2EC component or it is on a $y-z$-path in $H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ for $y,z\in\{u,v,w\}$, $y\neq z$ (hence Lemma 25 holds). In the remaining case $H_{2}^{\mathsf{C1}}$ is composed of three connected components, each containing exactly one vertex of $\{u,v,w\}$. Observe that each component is 2EC. Hence, Lemma 24 guarantees the existence of at least two edges $e_{1}$ and $e_{2}$ in $G_{2}$ between two pairs of connected components of $H_{2}^{\mathsf{C1}}$ such that $H_{2}^{\mathsf{C1}}\cup\{e_{1},e_{2}\}$ is connected. Thus, we can analogously to the previous case(s) show that $\mathsf{OPT}_{1}^{\mathsf{C1}}\cup H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ with $F^{\mathsf{C1}}\coloneq\{e_{1},e_{2}\}$ is a 2EC spanning subgraph of $G$, since in $H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ each edge is either in some 2EC component or it is on a $y-z$-path in $H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ for $y,z\in\{u,v,w\}$, $y\neq z$ (hence Lemma 25 holds). ∎
@@
If the condition in Algorithm 2 in Algorithm 2 applies and $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C2}}$, then there exist edges $F^{\mathsf{C2}}$ with $|F^{\mathsf{C2}}|\leq 1$ such that $H^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ and $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|\leq-2$.
@@
We first consider the case where $vw\in\mathsf{RED}(G_{2}^{\mathsf{C2}})$. Let $C_{1}^{\mathsf{C2}}(w)$ be the 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ that contains $w$. Since $C_{1}^{\mathsf{C2}}(w)$ is isolated in $\mathsf{OPT}_{1}^{\mathsf{C2}}$, Lemma 24 implies that there exists an edge $e$ in $G_{1}$ between $C_{1}^{\mathsf{C2}}(w)$ and $V(G_{1})\setminus C_{1}^{\mathsf{C2}}(w)$. Then, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ is connected and spanning for $G_{1}$, and every edge in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C2}}$. Similarly, every edge in $H_{2}^{\mathsf{C2}}$ is part of a 2EC component of $H_{2}^{\mathsf{C2}}$ or lies on a path in $H_{2}^{\mathsf{C2}}$ between any two distinct vertices of $\{u,v,w\}$ between which is a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$. Thus, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup H_{2}^{\mathsf{C2}}\cup F^{\mathsf{C2}}$ with $F^{\mathsf{C2}}\coloneq\{e\}$ is a 2EC spanning subgraph of $G$ by Lemma 25. Moreover, $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=1-3=-2$.
If $vw\notin\mathsf{RED}(G_{2}^{\mathsf{C2}})$, then every edge in $\mathsf{OPT}_{1}^{\mathsf{C2}}$ is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}$ between $u$ and $v$ between which is also a path in $H_{2}^{\mathsf{C2}}$. Further, every edge in $H_{2}^{\mathsf{C2}}$ is part of a 2EC component of $H_{2}^{\mathsf{C2}}$ or lies on a path in $H_{2}^{\mathsf{C2}}$ between $u$ and $v$, and there is a path between $u$ and $v$ in $\mathsf{OPT}_{1}^{\mathsf{C2}}$. Note that there cannot be an edge in $H_{2}^{\mathsf{C2}}$ that is on a path between $w$ and $u$ or $v$ but not in a 2EC component, because then $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ cannot be feasible for $G_{2}^{\mathsf{C2}}$ as $vw\notin\mathsf{RED}(G_{2}^{\mathsf{C2}})$. Thus, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup H_{2}^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ by Lemma 25. We have that $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=0-2=-2$. ∎
@@
If the condition in Algorithm 2 in Algorithm 2 applies and $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C2}}$, then there exist edges $F^{\mathsf{C2}}$ with $|F^{\mathsf{C2}}|\leq 1$ and $\mathsf{OPT}_{1}^{\mathsf{C2}}$ such that $H^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ and $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|\leq-3$.

### v2

@@ A.2.2 Auxiliary Lemmas for Algorithm˜3
Let $H\subseteq E$ be such that $H$ is connected and contains a spanning subgraph of $G$, and let $H_{i}\coloneq H\cap E(G_{i})$ for $i\in\{1,2\}$. If for all $i\in\{1,2\}$ every edge $e\in H_{i}$ is part of a 2EC component of $H_{i}$ or lies on an xyx-y path in $H_{i}$ where $x,y\in\{u,v,w\}$ such that $x\neq y$ and there exists an xyx-y path in $H_{j}$ where $j=1$ if $i=2$ and $j=2$ if $i=1$, then $H$ is a 2EC spanning subgraph of $G$.
@@
We show that every edge $e\in H$ is part of a 2EC component of $H$, which proves the lemma because $H$ is connected and spanning. If for all $i\in\{1,2\}$ every edge $e\in H_{i}$ is part of a 2EC component of $H_{i}$, it is also part of a 2EC component of $H$. Otherwise, for all $i\in\{1,2\}$, if $e$ lies on an xyx-y path in $H_{i}$ where $x,y\in\{u,v,w\}$ such that $x\neq y$ and there exists an xyx-y path in $H_{j}$ where $j=1$ if $i=2$ and $j=2$ if $i=1$, then $e$ is on a cycle of $H$ as both paths are edge-disjoint. ∎
@@
If the condition in Algorithm˜3 in Algorithm˜3 applies and both, $H^{\prime}_{1}$ and $H^{\prime}_{2}$ are 2EC spanning subgraphs of $G^{\prime}_{1}$ and $G^{\prime}_{2}$, respectively, then there exist edges $F^{\prime}$ with $|F^{\prime}|\leq 4$ such that $H^{\prime}$ is a 2EC spanning subgraph of $G$.
@@
Let $i\in\{1,2\}$, and let $H^{\prime}_{i}(u)$, $H^{\prime}_{i}(v)$, and $H^{\prime}_{i}(w)$ be the connected components of $H^{\prime}_{i}$ after decontracting $\{u,v,w\}$ that contain $u$, $v$ and $w$, respectively. By Lemma˜35, there exist edges $F_{i}\subseteq E(G_{i})$ such that $|F_{i}|\leq 2$ and $H^{\prime}_{i}\cup F_{i}$ is connected in $G_{i}$. Since $H^{\prime}_{i}$ is spanning in $G_{i}$, setting $F\coloneq F_{1}\cup F_{2}$ implies that $|F|\leq 4$ and that $H^{\prime}=H^{\prime}_{1}\cup H^{\prime}_{2}\cup F$ is spanning and connected in $G$. Moreover, observe that for $i\in\{1,2\}$ each edge $e\in H^{\prime}_{i}\cup F_{i}$ is either in a 2EC component of $H^{\prime}_{i}\cup F_{i}$ or lies on an xyx-y path in $H^{\prime}_{i}\cup F_{i}$, where $\in\{u,v,w\}, xyx\neq y$. Thus, Lemma˜36 implies that $H^{\prime}$ is a 2EC spanning subgraph of $G$. ∎
@@
If the condition in Algorithm˜3 in Algorithm˜3 applies and $\mathsf{RED}(G_{2}^{\mathsf{B1}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{B1}}$, then there exist edges $F^{\mathsf{B1}}$ with $|F^{\mathsf{B1}}|\leq 1$ such that $H^{\mathsf{B1}}$ is a 2EC spanning subgraph of $G$.
@@
Assume w.l.o.g. that $u$ and $v$ belong to the same 2EC component of $\mathsf{OPT}_{1}^{\mathsf{B1}}$. Let $H_{1}^{\mathsf{B1}}\coloneq\mathsf{OPT}_{1}^{\mathsf{B1}}$ and $C_{i}^{\mathsf{B1}}(x)$ be the 2EC component of $H_{i}^{\mathsf{B1}}$ that contains $x$, where $x\in\{u,v,w\}$ and $i\in\{1,2\}$. Since $C_{1}^{\mathsf{B1}}(u)=C_{1}^{\mathsf{B1}}(v)$, we have that in $\mathsf{OPT}_{1}^{\mathsf{B1}}\cup H_{2}^{\mathsf{B1}}$, the vertices of $C_{1}^{\mathsf{B1}}(u)$, $C_{2}^{\mathsf{B1}}(u)$, and $C_{2}^{\mathsf{B1}}(v)$ are in the same 2EC component. Furthermore, in $\mathsf{OPT}_{1}^{\mathsf{B1}}$, there is a connection between $C_{1}^{\mathsf{B1}}(u)$ and $C_{1}^{\mathsf{B1}}(w)$. If $w$ and $v$ are already connected in $H_{2}^{\mathsf{B1}}$ then we can set $F^{\mathsf{B1}}\coloneq\emptyset$ and see that this satisfies the lemma, since $H^{\mathsf{B1}}$ contains a spanning tree and every edge in $H^{\mathsf{B1}}$ is either in the 2EC component containing $u$ and $v$, in a 2EC component containing $w$ or on a wvw-v path in $G_{i}$, $i\in\{1,2\}$. Otherwise, if $w$ and $v$ are not in the same connected component in $H_{2}^{\mathsf{B1}}$, by Lemma˜35, there must be an edge $e$ from $C_{2}^{\mathsf{B1}}(w)$ to either $C_{2}^{\mathsf{B1}}(v)$ or $C_{2}^{\mathsf{B1}}(u)$ in $E(G_{2})\setminus H^{\mathsf{B1}}$. Now it can be easily verified that $F^{\mathsf{B1}}\coloneq\{e\}$ satisfies the lemma, since $H^{\mathsf{B1}}$ contains a spanning tree and every edge in $H^{\mathsf{B1}}$ is either in the 2EC component containing $u$ and $v$, in a 2EC component containing $w$ or on a wvw-v path in $G_{i}$, $i\in\{1,2\}$. ∎
@@
If the condition in Algorithm˜3 in Algorithm˜3 applies and $\mathsf{RED}(G_{2}^{\mathsf{B2}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{B2}}$, then there exist edges $F^{\mathsf{B2}}$ with $|F^{\mathsf{B2}}|\leq 2$ such that $H^{\mathsf{B2}}$ is a 2EC spanning subgraph of $G$.
@@
The proof is similar to the previous one. Assume w.l.o.g. that $u$ and $v$ belong to the same 2EC component of $\mathsf{OPT}_{1}^{\mathsf{B2}}$. Let $H_{1}^{\mathsf{B2}}\coloneq\mathsf{OPT}_{1}^{\mathsf{B2}}$ and $C_{i}^{\mathsf{B2}}(x)$ be the 2EC component of $H_{i}^{\mathsf{B2}}$ that contains $x$, where $x\in\{u,v,w\}$ and $i\in\{1,2\}$. Since $C_{1}^{\mathsf{B2}}(u)=C_{1}^{\mathsf{B2}}(v)$, we have that in $\mathsf{OPT}_{1}^{\mathsf{B2}}\cup H_{2}^{\mathsf{B2}}$, the vertices of $C_{1}^{\mathsf{B2}}(u)$, $C_{2}^{\mathsf{B2}}(u)$ and $C_{2}^{\mathsf{B2}}(v)$ are in the same 2EC component. Furthermore, in $\mathsf{OPT}_{1}^{\mathsf{B2}}$, there is no connection between $C_{1}^{\mathsf{B1}}(u)$ and $C_{1}^{\mathsf{B2}}(w)$, since it is of type $\mathsf{B2}$. Hence, by Lemma˜35, there must be an edge $e_{1}\in E(G_{1})\setminus\mathsf{OPT}_{1}^{\mathsf{B2}}$. Note that $\mathsf{OPT}_{1}^{\mathsf{B2}}\cup\{e_{1}\}$ is now a solution of type $\mathsf{B1}$ for $G_{1}$, and hence we can do the same as in the proof of the previous lemma and observe that $\lvert F^{\mathsf{B2}}\rvert\leq 2$. ∎
@@
If the condition in Algorithm˜3 in Algorithm˜3 applies and $\mathsf{RED}(G_{2}^{\mathsf{C1}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C1}}$, then there exist edges $F^{\mathsf{C1}}$ with $|F^{\mathsf{C1}}|\leq 2$ such that $H^{\mathsf{C1}}$ is a 2EC spanning subgraph of $G$.
@@
Let $C_{2}^{\mathsf{C1}}(x)$ be the 2EC component of $H_{2}^{\mathsf{C1}}$ that contains $x$, for $x\in\{u,v,w\}$. Note that $\mathsf{OPT}_{1}^{\mathsf{C1}}$ is connected and spanning for $G_{1}$ and that each edge $e\in\mathsf{OPT}_{1}^{\mathsf{C1}}$ is either in some 2EC component or it is on a yzy-z-path in $\mathsf{OPT}_{1}^{\mathsf{C1}}$ for $\in\{u,v,w\}, yzy\neq z$. We consider three cases. If $H_{2}^{\mathsf{C1}}$ is connected, then we claim that $F^{\mathsf{C1}}\coloneq\emptyset$ satisfies the lemma. In this case every edge of $H_{2}^{\mathsf{C1}}$ is part of a 2EC component of $H_{2}^{\mathsf{C1}}$, or it is on a yzy-z-path in $H_{2}^{\mathsf{C1}}$ for $\in\{u,v,w\}, yzy\neq z$. Hence, Lemma˜36 implies that $H_{2}^{\mathsf{C1}}\cup\mathsf{OPT}_{1}^{\mathsf{C1}}$ is a 2EC spanning subgraph of $G$.
Next, assume that $H_{2}^{\mathsf{C1}}$ is composed of two connected components, say w.l.o.g. one containing $u$ and $v$ and the other containing $w$, and observe that each component is 2EC. Then Lemma˜35 guarantees the existence of an edge $e$ in $G_{2}$ between $C_{2}^{\mathsf{C1}}(w)$ and $C_{2}^{\mathsf{C1}}(u)$ such that $H_{2}^{\mathsf{C1}}\cup\{e\}$ is connected. Thus, we can analogously to the previous case show that $\mathsf{OPT}_{1}^{\mathsf{C1}}\cup H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ with $F^{\mathsf{C1}}\coloneq\{e\}$ is a 2EC spanning subgraph of $G$, since in $H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ each edge is either in some 2EC component or it is on a yzy-z-path in $H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ for $\in\{u,v,w\}, yzy\neq z$ (hence Lemma˜36 holds). In the remaining case $H_{2}^{\mathsf{C1}}$ is composed of three connected components, each containing exactly one vertex of $\{u,v,w\}$. Observe that each component is 2EC. Hence, Lemma˜35 guarantees the existence of at least two edges $e_{1}$ and $e_{2}$ in $G_{2}$ between two pairs of connected components of $H_{2}^{\mathsf{C1}}$ such that $H_{2}^{\mathsf{C1}}\cup\{e_{1},e_{2}\}$ is connected. Thus, we can analogously to the previous cases show that $\mathsf{OPT}_{1}^{\mathsf{C1}}\cup H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ with $F^{\mathsf{C1}}\coloneq\{e_{1},e_{2}\}$ is a 2EC spanning subgraph of $G$, since in $H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ each edge is either in some 2EC component or it is on a yzy-z-path in $H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ for $\in\{u,v,w\}, yzy\neq z$ (hence Lemma˜36 holds). ∎
@@
If the condition in Algorithm˜3 in Algorithm˜3 applies and $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C2}}$, then there exist edges $F^{\mathsf{C2}}$ with $|F^{\mathsf{C2}}|\leq 1$ such that $H^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ and $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|\leq-2$.
@@
We first consider the case where $vw\in\mathsf{RED}(G_{2}^{\mathsf{C2}})$. Let $C_{1}^{\mathsf{C2}}(w)$ be the 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ that contains $w$. Since $C_{1}^{\mathsf{C2}}(w)$ is isolated in $\mathsf{OPT}_{1}^{\mathsf{C2}}$, Lemma˜35 implies that there exists an edge $e$ in $G_{1}$ between $C_{1}^{\mathsf{C2}}(w)$ and $V(G_{1})\setminus C_{1}^{\mathsf{C2}}(w)$. Then, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ is connected and spanning for $G_{1}$, and every edge in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C2}}$. Similarly, every edge in $H_{2}^{\mathsf{C2}}$ is part of a 2EC component of $H_{2}^{\mathsf{C2}}$ or lies on a path in $H_{2}^{\mathsf{C2}}$ between any two distinct vertices of $\{u,v,w\}$ between which is a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$. Thus, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup H_{2}^{\mathsf{C2}}\cup F^{\mathsf{C2}}$ with $F^{\mathsf{C2}}\coloneq\{e\}$ is a 2EC spanning subgraph of $G$ by Lemma˜36. Moreover, $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=1-3=-2$.
If $vw\notin\mathsf{RED}(G_{2}^{\mathsf{C2}})$, then every edge in $\mathsf{OPT}_{1}^{\mathsf{C2}}$ is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}$ between $u$ and $v$ between which is also a path in $H_{2}^{\mathsf{C2}}$. Further, every edge in $H_{2}^{\mathsf{C2}}$ is part of a 2EC component of $H_{2}^{\mathsf{C2}}$ or lies on a path in $H_{2}^{\mathsf{C2}}$ between $u$ and $v$, and there is a path between $u$ and $v$ in $\mathsf{OPT}_{1}^{\mathsf{C2}}$. Note that there cannot be an edge in $H_{2}^{\mathsf{C2}}$ that is on a path between $w$ and $u$ or $v$ but not in a 2EC component, because then $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ cannot be feasible for $G_{2}^{\mathsf{C2}}$ as $vw\notin\mathsf{RED}(G_{2}^{\mathsf{C2}})$. Thus, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup H_{2}^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ by Lemma˜36. We have that $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=0-2=-2$. ∎
@@
If the condition in Algorithm˜3 in Algorithm˜3 applies and $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C2}}$, then there exist edges $F^{\mathsf{C2}}$ with $|F^{\mathsf{C2}}|\leq 1$ and $\mathsf{OPT}_{1}^{\mathsf{C2}}$ such that $H^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ and $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|\leq-3$.